# On some dynamical features of the complete Moran model for neutral evolution in the presence of mutations

### 20 Feb 2024 v1

@@ 1 Introduction
It should be mentioned in this regard that the original paper by Moran [1] also took mutations into account and provided results for this case, see below for detail. Most of the literature dealing with the Moran mode (including that considering generalizations of the Moran model, see e.g. [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]), however, disregards the possibility of mutations; we will hence refer to the Moran model without mutations as the standard Moran model.
@@ 3 The Moran model in the presence of mutations
We will thus consider a simple Moran-like model in which mutations are also present; this will be dubbed the “complete” Moran model (as mentioned in the Introduction, this was actually considered by Moran in his seminal 1957 paper).
@@
    The chosen individuals reproduces, giving birth to a single descendant; this has the same allele of the parent with probability $(1-\mu)$, and the other allele with probability $\mu$ ;
@@
We stress that the process is quite obviously a Markov chain, and actually an ergodic one. Note that for $\mu=0$ we recover exactly the standard Moran model (which is not ergodic).
@@
We also note that the $F(x)$ with smaller $\mu$ will be more flat in the central region, and steeper at the borders, see Figure 1. They become of course singular in the limit $\mu\to 0$.
@@ 6.1 Dependence on the total population number
Refer to caption Refer to caption
$n=10$ $n=25$
Refer to caption Refer to caption
$n=50$ $n=100$
@@ 7.1 Dependence on the total population
Refer to caption Refer to caption
$n=10$ $n=25$
Refer to caption Refer to caption
$n=50$ $n=100$
@@ 7.2 Dependence on the mutation rate
Albeit in the complete Moran model fixation is not possible, and is substituted by reaching (temporarily) a “pure” state, the dependence of the mean hitting time on the mutation rate is specially interesting from the point of view of Evolutionary Dynamics (see the discussion in Sect. LABEL:sec:defix below).

### 23 Mar 2024 v2

# Open Communications in Nonlinear Mathematical Physics    ]ocnmp[   Vol.4 (2024) pp 1References   Article

@@ 1 Introduction
It should be mentioned in this regard that the original paper by Moran [1] also took mutations into account and provided results for this case, see below for detail. Most of the literature dealing with the Moran model (including that considering generalizations of the Moran model, see e.g. [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]), however, disregards the possibility of mutations; we will hence refer to the Moran model without mutations as the standard Moran model.
@@ 3 The Moran model in the presence of mutations
We will thus consider a simple Moran-like model in which mutations are also present; this will be dubbed the “complete” Moran model (as mentioned in the Introduction, this was actually considered by Moran in his seminal 1958 paper).
@@
    The chosen individuals reproduces, giving birth to a single descendant; this has the same allele of the parent with probability $(1-\mu)<1$, and the other allele with probability $\mu>0$ ;
@@
We stress that the process is quite obviously a Markov chain, and actually an ergodic one. Note that in the limit $\mu=0$ we recover exactly the standard Moran model (which is not ergodic, as the states $k=0$ and $k=N$ are absorbing).
@@
We also note that the functions $F(x)$ with smaller $\mu$ will be more flat in the central region, and steeper at the borders, see Figure 1. They become of course singular in the limit $\mu\to 0$.
@@ 6.1 Dependence on the total population number
Refer to caption Refer to caption $n=10$ $n=25$ Refer to caption Refer to caption $n=50$ $n=100$
@@ 7.1 Dependence on the total population
Refer to caption Refer to caption $n=10$ $n=25$ Refer to caption Refer to caption $n=50$ $n=100$
@@ 7.2 Dependence on the mutation rate
Albeit in the complete Moran model fixation is not possible, and is substituted by reaching (temporarily) a “pure” state, the dependence of the mean hitting time on the mutation rate is specially interesting from the point of view of Evolutionary Dynamics (see the discussion in Sect. 7.3 below).